OpenStudy (anonymous):

if y=sin^-1(5x), then dy/dx=

geerky42 (geerky42):

\(\Large \dfrac{d}{dx} sin^{-1} u = \dfrac{u'}{\sqrt{1-u^2}}\)

geerky42 (geerky42):

\(u = 5x\)

geerky42 (geerky42):

Is this clear? No?